# Sample-Efficient Constrained Reinforcement Learning with General Parameterization

### v1

@@ Appendix G Proof of Theorem 1
$\begin{split}\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{[}J_{r,\rho}^{\pi^{*}}&-J_{r,\rho}(\theta_{k})\bigg{]}+\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{[}-\lambda_{k}J_{c,\rho}(\theta_{k})\bigg{]}\\ &\leq\sqrt{\epsilon_{\mathrm{bias}}}+f_{0}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\dfrac{1}{H}+f_{1}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{3}}\dfrac{1}{K}+\dfrac{f_{2}}{(1-\gamma)^{2}}\zeta\end{split}$ (64)
G.1 Convergence Rate of the Objective Function
@@
$\begin{split}\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{[}J_{r,\rho}^{\pi^{*}}&-J_{r,\rho}(\theta_{k})\bigg{]}\leq\sqrt{\epsilon_{\mathrm{bias}}}+f_{0}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\dfrac{1}{H}\\ &+f_{1}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{3}}\dfrac{1}{K}+\dfrac{f_{2}}{(1-\gamma)^{2}}\zeta+\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{[}\lambda_{k}J_{c,\rho}(\theta_{k})\bigg{]}\end{split}$ (65)
@@
$\begin{split}\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{[}J_{r,\rho}^{\pi^{*}}&-J_{r,\rho}(\theta_{k})\bigg{]}\leq\sqrt{\epsilon_{\mathrm{bias}}}+f_{0}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\dfrac{1}{H}+f_{1}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{3}}\dfrac{1}{K}+\dfrac{f_{2}+1}{(1-\gamma)^{2}}\zeta\end{split}$ (70)
G.2 Rate of Constraint Violation
@@
$\begin{split}\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\left[(\lambda_{k}-\lambda_{\max})\hat{J}_{c,\rho}(\theta_{k})\right]&\overset{(a)}{\leq}\dfrac{|\lambda_{\max}|^{2}-|\lambda_{K}-\lambda_{\max}|^{2}}{2\zeta K}+\dfrac{\zeta}{2K}\sum_{k=0}^{K-1}\mathbf{E}\left[\hat{J}^{2}_{c,\rho}(\theta_{k})\right]\\ &\overset{(b)}{\leq}\dfrac{\lambda_{\max}^{2}}{2\zeta K}+\dfrac{\zeta}{(1-\gamma)^{2}}\end{split}$ (72)
where (a) uses $\lambda_{0}=0$ and (b) applies (69). Note that one can write the following using Lemma 1 and the observation that $\hat{J}_{c,\rho}(\theta_{k})$ and $\lambda_{k}$ are conditionally independent given $\theta_{k}$.
@@
$\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{[}(\lambda_{k}-\lambda_{\max})J_{c,\rho}(\theta_{k})\bigg{]}\leq\dfrac{\lambda_{\max}^{2}}{2\zeta K}+\dfrac{\zeta}{(1-\gamma)^{2}}$ (74)
@@
$\begin{split}\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{[}J_{r,\rho}^{\pi^{*}}&-J_{r,\rho}(\theta_{k})\bigg{]}+\lambda_{\max}\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{[}-J_{c,\rho}(\theta_{k})\bigg{]}\\ &\leq\sqrt{\epsilon_{\mathrm{bias}}}+f_{0}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\dfrac{1}{H}+f_{1}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{3}}\dfrac{1}{K}+\dfrac{f_{2}}{(1-\gamma)^{2}}\zeta+\dfrac{\lambda_{\max}^{2}}{2\zeta K}\end{split}$ (75)
@@
$\begin{split}J_{r,\rho}^{\pi^{*}}-\mathbf{E}\big{[}J_{r,\rho}^{\bar{\pi}}\big{]}+\lambda_{\max}\mathbf{E}\big{[}-J_{c,\rho}^{\bar{\pi}}\big{]}\leq\sqrt{\epsilon_{\mathrm{bias}}}&+f_{0}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\dfrac{1}{H}\\ &+f_{1}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{3}}\dfrac{1}{K}+\dfrac{f_{2}}{(1-\gamma)^{2}}\zeta+\dfrac{\lambda_{\max}^{2}}{2\zeta K}\end{split}$ (76)
@@
$\begin{split}\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\left[-J_{c,\rho}(\theta_{k})\right]\leq\dfrac{2\sqrt{\epsilon_{\mathrm{bias}}}}{\lambda_{\max}}&+2f_{0}\dfrac{(1+\lambda_{\max})}{\lambda_{\max}(1-\gamma)^{2}}\dfrac{1}{H}\\ &+2f_{1}\dfrac{(1+\lambda_{\max})}{\lambda_{\max}(1-\gamma)^{3}}\dfrac{1}{K}+\dfrac{2f_{2}}{\lambda_{\max}(1-\gamma)^{2}}\zeta+\dfrac{\lambda_{\max}}{\zeta K}\end{split}$ (77)
G.3 Final Result
Note from (77) that the terms related to $\zeta$ can be expressed as $\mathcal{O}(\zeta/(\lambda_{\max}(1-\gamma)^{2})+\lambda_{\max}/(\zeta K))$. Therefore, the optimal choice of $\zeta$ is $\zeta=\lambda_{\max}(1-\gamma)/\sqrt{K}$. Substituting $\lambda_{\max}=2/[(1-\gamma)c_{\mathrm{slater}}]\geq 2$ in (70), we get the rate of convergence of the objective as follows.
$\begin{split}\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{[}J_{r,\rho}^{\pi^{*}}&-J_{r,\rho}(\theta_{k})\bigg{]}\leq\sqrt{\epsilon_{\mathrm{bias}}}+\dfrac{3f_{0}c_{\mathrm{slater}}^{-1}}{(1-\gamma)^{3}}\dfrac{1}{H}+\dfrac{3f_{1}c_{\mathrm{slater}}^{-1}}{(1-\gamma)^{4}}\dfrac{1}{K}+\dfrac{2(f_{2}+1)c_{\mathrm{slater}}^{-1}}{(1-\gamma)^{2}\sqrt{K}}\end{split}$ (78)
@@ G.3 Final Result
$\begin{split}\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\left[-J_{c,\rho}(\theta_{k})\right]\leq(1-\gamma)c_{\mathrm{slater}}\sqrt{\epsilon_{\mathrm{bias}}}&+\dfrac{3f_{0}}{(1-\gamma)^{2}}\dfrac{1}{H}\\ &+\dfrac{3f_{1}}{(1-\gamma)^{3}}\dfrac{1}{K}+\dfrac{2f_{2}+1}{(1-\gamma)}\dfrac{1}{\sqrt{K}}\end{split}$ (79)
@@
$\begin{split}H&=\max\left\{\dfrac{9f_{0}c_{\mathrm{slater}}^{-1}}{(1-\gamma)^{3}}\epsilon^{-1},\dfrac{9f_{0}}{(1-\gamma)^{2}}\epsilon^{-1}\right\}=\mathcal{O}((1-\gamma)^{-3}\epsilon^{-1}),\\ K&=\max\left\{\dfrac{9f_{1}c_{\mathrm{slater}}^{-1}}{(1-\gamma)^{4}}\epsilon^{-1},\dfrac{9f_{1}}{(1-\gamma)^{3}}\epsilon^{-1},\dfrac{36(f_{2}+1)^{2}c_{\mathrm{slater}}^{-2}}{(1-\gamma)^{4}}\epsilon^{-2},\dfrac{9(2f_{2}+1)^{2}}{(1-\gamma)^{2}}\epsilon^{-2}\right\}\\ &=\mathcal{O}((1-\gamma)^{-4}\epsilon^{-2})\end{split}$ (80)
@@
Note that the expected number of steps required in Algorithm 19 is $\mathcal{O}((1-\gamma)^{-1})$. Therefore, the sample complexity required to ensure (81) is $\mathcal{O}((1-\gamma)^{-1}HK)=\mathcal{O}((1-\gamma)^{-8}\epsilon^{-3})$.
Appendix H Strong Duality and Related Lemmas
@@
[41] [Lemma 3] If $\lambda^{*}\triangleq{\arg\min}_{\lambda\geq 0}J_{\mathrm{D},\rho}^{\lambda}$ and $\pi^{*}$ is a solution of (2), then
@@
$\begin{split}\frac{C}{2}(-\tau)\leq(C-\lambda^{*})(-\tau)\leq J_{r,\rho}^{\pi^{*}}-J_{r,\rho}^{\bar{\pi}}+C(-\tau)\leq\zeta\end{split}$ (87)
This concludes the proof. ∎

### v3

@@ Appendix F Proof of Theorem 1
$\begin{split}\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{[}J_{r,\rho}^{\pi^{*}}&-J_{r,\rho}(\theta_{k})\bigg{]}+\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{[}-\lambda_{k}J_{c,\rho}(\theta_{k})\bigg{]}\\ &\leq\sqrt{\epsilon_{\mathrm{bias}}}+f_{0}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\exp\left(-\dfrac{\mu_{F}}{40G^{2}}H\right)+f_{1}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\dfrac{1}{\sqrt{K}}\end{split}$ (57)
F.1 Convergence Rate of the Objective Function
@@
$\begin{split}\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{[}J_{r,\rho}^{\pi^{*}}&-J_{r,\rho}(\theta_{k})\bigg{]}\leq\sqrt{\epsilon_{\mathrm{bias}}}+f_{0}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\exp\left(-\dfrac{\mu_{F}}{40G^{2}}H\right)\\ &+f_{1}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\dfrac{1}{\sqrt{K}}+\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{[}\lambda_{k}J_{c,\rho}(\theta_{k})\bigg{]}\end{split}$ (58)
@@
$\begin{split}&\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{[}J_{r,\rho}^{\pi^{*}}-J_{r,\rho}(\theta_{k})\bigg{]}\\ &\leq\sqrt{\epsilon_{\mathrm{bias}}}+f_{0}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\exp\left(-\dfrac{\mu_{F}}{40G^{2}}H\right)+f_{1}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\dfrac{1}{\sqrt{K}}+\dfrac{\zeta}{(1-\gamma)^{2}}\\ &\overset{(a)}{=}\sqrt{\epsilon_{\mathrm{bias}}}+f_{0}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\exp\left(-\dfrac{\mu_{F}}{40G^{2}}H\right)+f_{1}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\dfrac{1}{\sqrt{K}}+\dfrac{\lambda_{\max}}{(1-\gamma)}\dfrac{1}{\sqrt{K}}\\ &\leq\sqrt{\epsilon_{\mathrm{bias}}}+f_{0}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\exp\left(-\dfrac{\mu_{F}}{40G^{2}}H\right)+(f_{1}+1)\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\dfrac{1}{\sqrt{K}}\end{split}$ (63)
where $(a)$ uses the substitution $\zeta=\lambda_{\max}(1-\gamma)/\sqrt{K}$.
F.2 Rate of Constraint Violation
@@
$\begin{split}\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\left[(\lambda_{k}-\lambda_{\max})\hat{J}_{c,\rho}(\theta_{k})\right]&\overset{(a)}{\leq}\dfrac{|\lambda_{\max}|^{2}-|\lambda_{K}-\lambda_{\max}|^{2}}{2\zeta K}+\dfrac{\zeta}{2K}\sum_{k=0}^{K-1}\mathbf{E}\left[\hat{J}^{2}_{c,\rho}(\theta_{k})\right]\\ &\overset{(b)}{\leq}\dfrac{\lambda_{\max}^{2}}{2\zeta K}+\dfrac{\zeta}{(1-\gamma)^{2}}\overset{(c)}{=}\dfrac{3\lambda_{\max}}{2(1-\gamma)}\dfrac{1}{\sqrt{K}}\end{split}$ (65)
where $(a)$ utilises $\lambda_{0}=0$, $(b)$ applies (62), and $(c)$ is derived using $\zeta=\lambda_{\max}(1-\gamma)/\sqrt{K}$. Note that one can write the following using Lemma 1 and the observation that $\hat{J}_{c,\rho}(\theta_{k})$ and $\lambda_{k}$ are conditionally independent given $\theta_{k}$.
@@
$\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{[}(\lambda_{k}-\lambda_{\max})J_{c,\rho}(\theta_{k})\bigg{]}\leq\dfrac{3\lambda_{\max}}{2(1-\gamma)}\dfrac{1}{\sqrt{K}}\leq\dfrac{2(1+\lambda_{\max})}{(1-\gamma)^{2}\sqrt{K}}$ (67)
@@
$\begin{split}\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{[}J_{r,\rho}^{\pi^{*}}&-J_{r,\rho}(\theta_{k})\bigg{]}+\lambda_{\max}\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{[}-J_{c,\rho}(\theta_{k})\bigg{]}\\ &\leq\sqrt{\epsilon_{\mathrm{bias}}}+f_{0}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\exp\left(-\dfrac{\mu_{F}}{40G^{2}}H\right)+(f_{1}+2)\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\dfrac{1}{\sqrt{K}}\end{split}$ (68)
@@
$\begin{split}J_{r,\rho}^{\pi^{*}}-\mathbf{E}\big{[}J_{r,\rho}^{\bar{\pi}}\big{]}+\lambda_{\max}\mathbf{E}\big{[}-J_{c,\rho}^{\bar{\pi}}\big{]}\leq\sqrt{\epsilon_{\mathrm{bias}}}&+f_{0}\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\exp\left(-\dfrac{\mu_{F}}{40G^{2}}H\right)\\ &+(f_{1}+2)\dfrac{(1+\lambda_{\max})}{(1-\gamma)^{2}}\dfrac{1}{\sqrt{K}}\end{split}$ (69)
@@
$\begin{split}\dfrac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\left[-J_{c,\rho}(\theta_{k})\right]\leq\dfrac{2\sqrt{\epsilon_{\mathrm{bias}}}}{\lambda_{\max}}&+2f_{0}\dfrac{(1+\lambda_{\max})}{\lambda_{\max}(1-\gamma)^{2}}\exp\left(-\dfrac{\mu_{F}}{40G^{2}}H\right)\\ &+2(f_{1}+2)\dfrac{(1+\lambda_{\max})}{\lambda_{\max}(1-\gamma)^{2}}\dfrac{1}{\sqrt{K}}\end{split}$ (70)
F.3 Final Result
Substituting $\lambda_{\max}=2/[(1-\gamma)c_{\mathrm{slater}}]\geq 2$ in (63), we get the rate of convergence of the objective as follows.
$\begin{split}\frac{1}{K}\sum_{k=0}^{K-1}\mathbf{E}\bigg{[}J_{r,\rho}^{\pi^{*}}&-J_{r,\rho}(\theta_{k})\bigg{]}\\ &\leq\sqrt{\epsilon_{\mathrm{bias}}}+\dfrac{3f_{0}c_{\mathrm{slater}}^{-1}}{(1-\gamma)^{3}}\exp\left(-\dfrac{\mu_{F}}{40G^{2}}H\right)+\dfrac{3(f_{1}+1)c_{\mathrm{slater}}^{-1}}{(1-\gamma)^{3}}\dfrac{1}{\sqrt{K}}\end{split}$ (71)
@@ F.3 Final Result
$\begin{split}\dfrac{1}{K}\sum_{k=0}^{K-1}&\mathbf{E}\left[-J_{c,\rho}(\theta_{k})\right]\\ &\leq(1-\gamma)c_{\mathrm{slater}}\sqrt{\epsilon_{\mathrm{bias}}}+\dfrac{3f_{0}}{(1-\gamma)^{2}}\exp\left(-\dfrac{\mu_{F}}{40G^{2}}H\right)+\dfrac{3(f_{1}+2)}{(1-\gamma)^{2}}\dfrac{1}{\sqrt{K}}\end{split}$ (72)
@@
$\begin{split}H&=\dfrac{40G^{2}}{\mu_{F}}\log\left(\max\left\{\dfrac{6f_{0}c_{\mathrm{slater}}^{-1}}{(1-\gamma)^{3}}\epsilon^{-1},\dfrac{6f_{0}}{(1-\gamma)^{2}}\epsilon^{-1}\right\}\right)=\mathcal{O}(\log(\epsilon^{-1})),\\ K&=\max\left\{\dfrac{36(f_{1}+1)^{2}c_{\mathrm{slater}}^{-2}}{(1-\gamma)^{6}}\epsilon^{-2},\dfrac{36(f_{1}+2)^{2}}{(1-\gamma)^{4}}\epsilon^{-2}\right\}=\mathcal{O}((1-\gamma)^{-6}\epsilon^{-2})\end{split}$ (73)
@@
Note that the expected number of steps required in Algorithm 19 is $\mathcal{O}((1-\gamma)^{-1})$. Therefore, the sample complexity required to ensure (74) is $\mathcal{O}((1-\gamma)^{-1}HK)=\tilde{\mathcal{O}}((1-\gamma)^{-7}\epsilon^{-2})$. Finally, note that Lemma 5 requires $H>\bar{C}\frac{G^{2}}{\mu_{F}}\log\left(\sqrt{\mathrm{d}}\frac{G^{2}}{\mu_{F}}\right)$. This can be ensured if $\epsilon$ is sufficiently small.
Appendix G Strong Duality and Related Lemmas
@@
[40] [Lemma 3] If $\lambda^{*}\triangleq{\arg\min}_{\lambda\geq 0}J_{\mathrm{D},\rho}^{\lambda}$ and $\pi^{*}$ is a solution of (2), then the following holds whenever Assumption 1 is true.
@@
$\begin{split}(C-\lambda^{*})(-\tau)\leq J_{r,\rho}^{\pi^{*}}-J_{r,\rho}^{\bar{\pi}}+C(-\tau)\leq\zeta\end{split}$ (80)
This leads to $-\tau\leq\zeta/(C-\lambda^{*})\leq 2\zeta/C$. ∎